A boat covers 30km upstream and 22 km downstrean in 8 hours. Also,it takes 4 more hours to travel 40 km upstream and 44 km downstrean. Find the speed of the boat in still water and that of the stream.
Answer by dkppathak(439)

let speed of boat in still water is x km /h
and speed of current of water is y km/h
speed up stream =x-y km/h
speed of down stream =x+y km/h
speed =distance / time
or
time =distance/ speed
t1=30/x-y
t2 =22/x+y
as per given conditions
t1 +t2 =8
30/x-y +22/x+y =8 (i)
similarly
40/x-y +44/x+y =12 (ii)
let x-y =u
x+y =v
we can rewrite equation as
30/u +22/v = 8 (iii)
40/u +44/v = 12 (iv)
by solving equations by elimination method
we can solve U =5
and V=11
by substitute the value of u and v
we can rewrite equations as
x-y =5
x +y =11
by solving these equations
we can get
x=8 km/h
y=3 km/h
ANSWER speed of boat in still water is 8km/h
and speed of current of water is 3 km/h